# Topological Phases, Criticality, and Mixed State Order in a Hubbard Quantum Simulator

### v1

@@ Acknowledgements
The experimental team is grateful for the contributions to building the experiment from A. Hebert, A. Krahn, G. Phelps, R. Groth, S. F. Ozturk, S. Ebadi, S. Dickerson, F. Ferlaino, V. Singh, and V. Kaxiras. We acknowledge fruitful discussions with D. Mark and Y. Wang, and we thank Ruochen Ma for insightful comments on the manuscript. We are supported by U.S. Department of Energy Quantum Systems Accelerator DE-AC02-05CH11231, National Science Foundation Center for Ultracold Atoms PHY-1734011, Army Research Office Defense University Research Instrumentation Program W911NF2010104, Office of Naval Research Vannevar Bush Faculty Fellowship N00014-18-1-2863, Gordon and Betty Moore Foundation Grant GBMF11521, and Defense Advanced Research Projects Agency Optimization with Noisy Intermediate-Scale Quantum devices W911NF-20-1-0021. A.D. acknowledges support from the NSF Graduate Research Fellowship Program (grant DGE2140743). R.S. acknowledges support from the Department of Energy Computational Science Graduate Fellowship (CSGF) under Award Number DE-SC0022158. C.D. was supported with the ITAMP grant No. 2116679.

### v2

@@ VI Conclusion and outlook
Connection between the MI-HI transition and the CSPT transition
Let us recall that, in the absence of the staggered potential ($\mu_{\text{S}}=0$), the phase diagram of the dipolar Bose-Hubbard model is theoretically predicted to contain at least three insulating phases as a function of the on-site interaction strength $U$ and the characteristic strength of the long-range interactions $V$ (schematically reproduced in main text Fig. 5a) [1, 2]. Two of these are the Mott insulator (MI) and the charge density wave (CDW), the latter of which spontaneously breaks the translation symmetry $\mathbb{Z}$ of the model at $\mu_{\text{S}}=0$ down to $2\mathbb{Z}$. The last phase is the Haldane Insulator (HI)—an entangled SPT phase that is separated from the Mott insulator by a quantum phase transition in the presence of the crystalline bond-centered inversion symmetry and parity symmetry of the dipolar Bose-Hubbard model at $\mu_{\text{S}}=0$ [7]. In this sense, the HI is a CSPT phase with protected quantum entanglement.
To see how the Haldane to Mott insulator transition is connected to the transition between CSPT-1 and CSPT-2 [38], let us note that under adding a staggered potential, the Mott insulator morphs into CSPT-1 manifestly. To gain an intuitive understanding of the relationship between HI and CSPT-2, we use the “dressed quasi-particle” picture of the CSPTs introduced in the main text (Section II).
In particular, the HI can be thought of as a case where bosons in our system are each delocalized across the bonds of the lattice, as shown in the main text Fig. 5a—a configuration that is symmetric under the parity, site- and bond-centered inversion symmetries of the model. Concretely, a model wavefunction for this phase is given by,
which captures the bond delocalization. The string order parameter in the main text Eq. (5) is similar to the parity string operator we used for CSPT-1 and CSPT-2; however, now there is an endpoint dressing $\delta\hat{n}=\hat{n}-1$. The reason for this is that without the dressing, the expected parity would vanish since the bond-delocalized bosons are equally likely to be inside as outside the region of interest. The dressing accounts for this, and the nonzero value of $\mathcal{S}_{x_{0}}(r)$ is predicted to be positive (for odd lengths), precisely consistent with the phenomenology of CSPT-2 and revealing their connection. Note that the endpoint operator $\delta n$ transforms to $-\delta n$ under the approximate charge-conjugation symmetry of the chain—it is precisely the transformation law of this endpoint decoration that serves as a topological invariant of the HI [10, 11, 12].
Approximate charge conjugation symmetry
@@ Acknowledgements
The experimental team is grateful for the contributions to building the experiment from A. Hebert, A. Krahn, G. Phelps, R. Groth, S. F. Ozturk, S. Ebadi, S. Dickerson, F. Ferlaino, V. Singh, and V. Kaxiras. We acknowledge fruitful discussions with Hans Peter Büchler, Daniel Mark, Lode Pollet and Yuxin Wang, and we thank Ruochen Ma for insightful comments on the manuscript. We are supported by U.S. Department of Energy Quantum Systems Accelerator DE-AC02-05CH11231, National Science Foundation Center for Ultracold Atoms PHY-1734011, Army Research Office Defense University Research Instrumentation Program W911NF2010104, Office of Naval Research Vannevar Bush Faculty Fellowship N00014-18-1-2863, Gordon and Betty Moore Foundation Grant GBMF11521, and Defense Advanced Research Projects Agency Optimization with Noisy Intermediate-Scale Quantum devices W911NF-20-1-0021. A.D. acknowledges support from the NSF Graduate Research Fellowship Program (grant DGE2140743). R.S. acknowledges support from the Department of Energy Computational Science Graduate Fellowship (CSGF) under Award Number DE-SC0022158. C.D. was supported with the ITAMP grant No. 2116679.